OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? Answer y - 4x = -3 y = 4x - 3 -4x = -y - 3 x = (y + 3)

OpenStudy (anonymous):

y = 4x - 3

OpenStudy (anonymous):

the bottom one supposed to have 1 over 5 after the equal sighn

OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation y - 4x = -3 Answer slope = -1; y-intercept = (0, -3) slope = - ; y-intercept = (0, 3) slope = 4; y-intercept = (0, -3) slope = 4; y-intercept = (0, 3)

OpenStudy (anonymous):

second ones 1 over 4